### v4

# A Practical Examination of AI-Generated Text Detectors for Large Language Models

Abstract
The proliferation of large language models has raised growing concerns about their misuse, particularly in cases where AI-generated text is falsely attributed to human authors. Machine-generated content detectors claim to effectively identify such text under various conditions and from any language model. This paper critically evaluates these claims by assessing several popular detectors (RADAR, Wild, T5Sentinel, Fast-DetectGPT, PHD, LogRank, Binoculars) on a range of domains, datasets, and models that these detectors have not previously encountered. We employ various prompting strategies to simulate practical adversarial attacks, demonstrating that even moderate efforts can significantly evade detection. We emphasize the importance of the true positive rate at a specific false positive rate (TPR@FPR) metric and demonstrate that these detectors perform poorly in certain settings, with TPR@.01 as low as 0%. Our findings suggest that both trained and zero-shot detectors struggle to maintain high sensitivity while achieving a reasonable true positive rate. All code and data necessary to reproduce our experiments are available at https://github.com/LeiLiLab/llm-detector-eval.
A Practical Examination of AI-Generated Text Detectors for Large Language Models
Brian Tufts Carnegie Mellon University btufts@cs.cmu.edu                        Xuandong Zhao UC Berkeley xuandongzhao@berkeley.edu                        Lei Li Carnegie Mellon University leili@cs.cmu.edu
1 Introduction
Large language models (LLMs) are becoming increasingly accessible and powerful, leading to numerous beneficial applications (Touvron et al., 2023; Achiam et al., 2023). However, they also pose risks if used maliciously, such as generating fake news articles, facilitating academic plagiarism or spam content (Feng et al., 2024; Zellers et al., 2019b; Perkins, 2023; Fraser et al., 2024). The potential for misuse of LLMs has become a significant concern for major tech corporations, particularly in light of the 2024 elections in the united states. At the Munich Security Conference on February 16th, 2024, these companies pledged to combat misleading machine-generated content, acknowledging the potential of AI to deceptively influence electoral outcomes (Accord, 2024). As a result, there is a growing need to develop reliable methods for differentiating between LLM-generated and human-written content. To ensure the effectiveness and accountability of LLM detection methods, continuous evaluation of popular techniques is crucial.
Many methods have been released recently that claim to have a strong ability to detect the difference between AI-generated and human-generated texts. These detectors primarily fall into three categories: trained detectors, zero-shot detectors, and watermarking techniques (Yang et al., 2023b; Ghosal et al., 2023; Tang et al., 2023). Trained detectors utilize datasets of human and AI-generated texts and train a binary classification model to detect the source of a text (Zellers et al., 2019b; Hovy, 2016; Hu et al., 2023; Tian and Cui, 2023; Verma et al., 2024). Zero-shot detection utilizes a language model’s inherent traits to identify text it generates, without explicit training for detection tasks other than calibrating a threshold for detection in some cases (Gehrmann et al., 2019; Mitchell et al., 2023; Bao et al., 2024; Yang et al., 2023a; Venkatraman et al., 2024). Watermarking is another technique in which the model owner embeds a specific probabilistic pattern into the text to make it detectable Kirchenbauer et al. (2023). However, watermarking requires the model owner to add the signal, and its design has theoretical guarantees; we do not evaluate watermarking models in this study.
In this paper, we test the robustness of these detection methods to unseen models, data sources, and adversarial prompting. To do this, we treat all model-generated text as a black box generation. That is, none of the detectors know the source of the text or have access to the model generating the text. This presents the most realistic scenario where the user is presented with text and wants to know if it is AI-generated or not. Our contributions can be summarized as follows:
  • We conduct a thorough evaluation of AI-generated text detectors on unseen models and tasks, providing insights into their effectiveness in real-world settings.
  • We analyze the performance of various detectors under practical adversarial prompting, exploring the extent to which prompting can be used to evade detection.
  • We demonstrate that high AUROC scores, which are often used as a measure of performance in classification tasks, do not necessarily translate to practical usage for machine-generated text detection. Instead, we motivate using the metric of true positive rate (TPR) at a 1% false positive rate (FPR) threshold as a more reliable indicator of a detector’s effectiveness in practice.
2 Related Work and Background
There is a variety of related work that discusses text detectors. These works cover different aspects, such as the text detectors themselves, their types, evaluation, and red-teaming of detectors.
Text Detectors.
Machine-generated text detectors can be divided into trained classifiers, zero-shot classifiers, and watermark methods (Yang et al., 2023b; Hans et al., 2024; Ghosal et al., 2023; Jawahar et al., 2020). (1) Trained detectors use classification models to determine if the text is machine-generated or human-written (Zellers et al., 2019b; Hovy, 2016; Hu et al., 2023; Tian and Cui, 2023; Verma et al., 2024). However, the increasing prevalence of machine-generated content (European-Union, 2022) makes it difficult to label human-generated work for training, as even humans find it hard to distinguish between the two (Darda et al., 2023). (2) Zero-shot detectors leverage intrinsic statistical differences between machine-generated and human-generated text (Gehrmann et al., 2019; Mitchell et al., 2023; Bao et al., 2024; Yang et al., 2023a; Venkatraman et al., 2024). Proposed methods include using entropy (Lavergne et al., 2008), log probability (Solaiman et al., 2019), and more recently, intrinsic dimensionality (Tulchinskii et al., 2023). (3) Watermark-based detection, introduced by Kirchenbauer et al. (2023), involves embedding a hidden but detectable pattern in the generated output. Various enhancements to this method have been suggested (e.g., Zhao et al. (2023); Lee et al. (2023)). This paper focuses on the black-box setting, which closely resembles real-world detection scenarios. Watermarking is not tested due to its guaranteed detectability and low false positive rates (e.g., (Zhao et al., 2023)). The primary concern is detecting un-watermarked text, as it is the most commonly encountered and poses the greatest threat.
Evaluation of Text Detectors.
The most commonly utilized metric in evaluating detectors is the area under the receiver operating curve (AUROC) (Mitchell et al., 2023; Sadasivan et al., 2023). Although it offers a reasonable estimate of detector performance, research by Krishna et al. (2023); Yang et al. (2023a), and our experimental results demonstrate that there can be a substantial difference in performance between two models with AUROC values nearing the maximum of 1.0. Consequently, the true positive rate at a fixed false positive rate (TPR@FPR) presents a more accurate representation of a detector’s practical effectiveness. Both AUROC and true positive rate at a fixed false positive are important metrics for a complete evaluation of text detectors.
Redteaming Language Model Detectors.
AI text detectors are increasingly evaluated in red teaming scenarios, with recent contributions from Zhu et al. (2023); Chakraborty et al. (2023); Kumarage et al. (2023); Shi et al. (2024); Wang et al. (2024). Shi et al. (2024) identifies two main evasion techniques: word substitution and instructional prompts. Word substitution includes query-based methods, which iteratively select low detection score substitutions, and query-free methods, which use random substitutions. Instructional prompts, akin to jailbreaking, instruct the model to mimic a human-written sample. Query-based word substitution proved most effective, reducing the True Positive Rate (TPR) to less than 5% at a 40% False Positive Rate (FPR) against DetectGPT. Wang et al. (2024) explore robustness testing of language model detectors with three editing attacks: typo insertion, homoglyph alteration, and format character editing. Typo insertion adds typos, homoglyph alteration replaces characters with similar shapes, and format character editing uses invisible text disruptions. Paraphrasing attacks, noted by Krishna et al. (2023), include synonym substitution (model-free and model-assisted), span perturbations (masking and refilling random spans), and paraphrasing at sentence and text levels.
Method Datasets
RADAR OpenWebText Corpus (Gokaslan et al., 2019), Xsum (Narayan et al., 2018), SQuAD (Rajpurkar et al., 2016), Reddit Writing Prompts (Fan et al., 2018), and TOEFL (Liang et al., 2023)
Wild Reddit CMV sub-community comments (Tan et al., 2016), Yelp Reviews (Zhang et al., 2015), Xsum (Narayan et al., 2018), TLDR_news [^1] https://huggingface.co/datasets/JulesBelveze/tldr_news, ELI5 dataset (Fan et al., 2019), Reddit Writing Prompts (Fan et al., 2018), ROCStories Corpora (Mostafazadeh et al., 2016), HellaSwag (Zellers et al., 2019a), SQuAD (Rajpurkar et al., 2016), and SciGen (Moosavi et al., 2021)
T5Sentinel OpenWebText Corpus (Gokaslan et al., 2019)
Fast-DetectGPT Xsum (Narayan et al., 2018), SQuAD (Rajpurkar et al., 2016), Reddit Writing Prompts (Fan et al., 2018), WMT16 English and German (Bojar et al., 2017), PubMedQA (Jin et al., 2019)
PHD Wiki40b (Guo et al., 2020), Reddit Writing Prompts (Fan et al., 2018), WikiM (Krishna et al., 2023), StackExchange (Tulchinskii et al., 2023)
LogRank Xsum (Narayan et al., 2018), SQuAD (Rajpurkar et al., 2016), Reddit Writing Prompts (Fan et al., 2018)
Binoculars CCNews (Hamborg et al., 2017), PubMed (Sen et al., 2008), CNN (Hermann et al., 2015), ORCA (Lian et al., 2023)
Table 1: Datasets used for training and evaluation by each model. To avoid data leakage and cherry-picking, these datasets are excluded from the current study.
Evaluated Detectors and Datasets.
In our paper, we evaluate seven representative detectors: RADAR (Hu et al., 2023), Detection in the Wild (Wild) (Li et al., 2024), T5Sentinel (Chen et al., 2023), Fast-DetectGPT (Bao et al., 2024), PHD (Tulchinskii et al., 2023), LogRank (Ippolito et al., 2020) [^2] LogRank has been evaluated on many datasets, we report the ones from Mitchell et al. (2023)., and Binoculars (Hans et al., 2024). RADAR, Wild, and T5Sentinel are trained detectors, while Fast-DetectGPT, PHD, LogRank, and Binoculars are zero-shot detectors. To ensure a fair comparison and assess the detectors’ ability to generalize to new data, we carefully select datasets that have not been used in the training or evaluation of these detectors. Table 1 presents an overview of the datasets and domains on which each detector has been evaluated. Several datasets, such as Xsum, SQuAD, and Reddit Writing Prompts, have been used in the evaluation or training of multiple detectors. Although these detectors achieve strong Area Under the Receiver Operating Characteristic (AUROC) scores on these datasets, they do not report the True Positive Rate at a set False Positive Rate (TPR@FPR), which is a crucial metric in real-world scenarios. To address this gap, we aim to evaluate all seven detectors on the same datasets using both AUROC and TPR at FPR metrics.
Comparison to Previous Works.
There are some other papers that have explored similar work to ours, specifically Wang et al. (2024) and Dugan et al. (2024). Our work differs from theirs in some important ways. We do not focus as much on the various methods of red-teaming the detectors in complicated ways. Rather, we explore some more natural methods that an average person might utilize in practice. We also explore in more depth the variability in detector capabilities across various tasks and languages with discussion on potential sources of that difference. And lastly, we utilize newer models, which gives insight into the adaptability of the detectors.
3 Benchmarking Procedure
Our benchmarking method involves compiling datasets that have not been encountered by any of the detectors during their training or evaluation phases. This approach ensures that the datasets represent new, unseen data and prevents the possibility of data contamination. For zero-shot detectors, this methodology eliminates the risk of using cherry-picked datasets that may bias the evaluation. For trained detectors, this reduces the risk of data leakage and tests on out-of-domain data. Furthermore, we assess the model’s performance across a diverse range of domains that the detectors may not have been previously evaluated against. This comprehensive evaluation strategy allows for a more robust assessment of the detectors’ generalization capabilities. Additionally, we evaluate the detectors on a variety of language models that they have not encountered before. This enables us to examine the detectors’ performance on unfamiliar language models, providing a more comprehensive understanding of their effectiveness and adaptability.
3.1 Datasets
We evaluate each of the detectors on seven different tasks with three of the tasks, question answering, summarization, and dialogue writing, including multilingual results. The datasets chosen for each domain are as follows:
  • Question Answering: The MFAQ dataset (De Bruyn et al., 2021) was used for this domain. It contains over one million question-answer pairs in various languages. We used the English, Spanish, French, and Chinese subsets.
  • Summarization: We used the MTG summarization dataset (Chen et al., 2022) for this task. The complete multilingual dataset comprises roughly 200k summarizations. We utilized the English, Spanish, French, and Chinese subsets.
  • Dialogue Writing: For this task, we utilized the MSAMSum dataset, a translated version of the SAMSum dataset(Feng et al., 2022; Gliwa et al., 2019). This dataset consists of over 16k dialogues with summaries in six languages. We utilized English, Spanish, French, and Chinese for consistency with the other multilingual domains.
  • Code: We used the APPS dataset (Hendrycks et al., 2021), which contains 10k code questions and solutions. The subset used was randomly selected from all the data included in APPS.
  • Abstract Writing: For this task, we utilized the Arxiv section of the scientific papers dataset (Cohan et al., 2018) to avoid potential bias, as some detectors have previously been exposed to PubMed data. Additionally, we only selected papers published in 2020 or earlier to remove potential LLM influence.
  • Review Writing: The PeerRead dataset was used for the review writing task (Kang et al., 2018). PeerRead contains over 10k peer reviews written by experts corresponding to the paper that they were written for.
  • Translation: We used the Par3 dataset (Karpinska et al., 2022), which provides paragraph level translations from public-domain foreign language novels. Each paragraph includes at least 2 human translations of which we selected only one to represent human translation.
3.2 Large Language Models
Our objective is to evaluate the detectors on models that they have not previously been trained or assessed on to gauge their generalization capabilities. We evaluated 4 different models across every task. The models we use are Llama-3-Instruct 8B (AI@Meta, 2024), Mistral-Instruct-v0.3 (Jiang et al., 2023), Phi-3-Mini-Instruct 4k (Abdin et al., 2024), and GPT-4o.
3.3 Detection Models
The detection models were chosen from the newest and highest performing detectors in their respective categories. Our goal was to represent both trained and zero-shot detectors. As previously mentioned, the trained detectors we are using are RADAR (Hu et al., 2023), Detection in the Wild (Wild) (Li et al., 2024), and T5Sentinel (Chen et al., 2023). The zero-shot detectors we are using are Fast-DetectGPT (Bao et al., 2024), GPTID (Tulchinskii et al., 2023), LogRank (Ippolito et al., 2020), and Binoculars (Hans et al., 2024). Each of the zero-shot detectors utilize a generating model as a part of their detection process. We utilize the same underlying models as reported by each respective zero-shot model’s original publication listed in Table 2. We also evaluate every zero-shot method using three of the other underlying models for a more accurate comparison. This is notably unfair to the Binoculars method, which uses two different underlying models: base and instruction tuned. We replace both with the same model for these experiments because not all models have both base and instruction tuned versions.
Notably, we did not include any watermark detectors. The primary reason for this is that the evaluation techniques we use over various models would not work with watermark detection. While watermark detection has shown strong performance (Kirchenbauer et al., 2023), they have a significant drawback in that they only work if a model applies a watermark. In this paper, we assume a scenario in which no watermark is applied or it is unknown whether a watermark is applied. Therefore, we must turn to other detection methods.
Method Model
Fast-DetectGPT GPT-Neo-2.7B (Black et al., 2021)
GPTID Roberta-Base Liu et al. (2019)
LogRank GPT2-Medium (Radford et al., 2019)
Binoculars Falcon-7B, Falcon-7B-Instruct (Almazrouei et al., 2023)
Table 2: Underlying models utilized by each zero-shot detection method.
3.4 Evaluation Metrics
In this study, we evaluate machine-generated text detectors using AUROC and TPR at a fixed FPR. Our findings, consistent with prior research (Krishna et al., 2023; Yang et al., 2023a), suggest that AUROC alone may not reflect a detector’s practical effectiveness, as a high AUROC score can still correspond to significant false positive rates. This is critical since false positives, particularly in fields like academia and media, can have severe consequences. We argue that TPR at a given FPR should be the standard evaluation metric, as demonstrated by a detector achieving a 0.89 AUROC but less than 20% TPR at a 1% FPR on a task.
3.5 Red Teaming
We employ two different methods of prompting for every task: plain prompting and adversarial prompting. Plain prompting involves using a typical assistant system prompt and providing the model with the same input that was given to the human for human-generated content. Adversarial prompting, on the other hand, requests that the model try to act more like a person. Examples of the question answering plain and adversarial prompts [^3] The others can be found in the appendix Table 13. are shown as follows:
Plain Prompt Example: Question Answering You are a helpful question answering assistant that will answer a single question as completely as possible given the information in the question. Do NOT use any markdown, bullet, or numbered list formatting. The assistant will use ONLY paragraph formatting. **Respond only in {language}**.
Adversarial Prompt Example: Question Answering {Question answering prompt} Try to sound as human as possible.
We also conducted experiments using the LLMs as writing assistants. Specifically, we requested that the model rewrite the human response and improve upon its clarity and professionalism. This represents a scenario where a person will write down an answer first and then request that a model make their answer better before presenting it. The specific prompt we used it as follow:
Rewriting Prompt You are a helpful writing assistant. Rewrite the following text to improve clarity and professionalism. Do not provide any other text. Only provide the rewritten text.
4 Experiment
4.1 Dataset Processing
Each dataset undergoes additional processing to prepare it for detection tasks. Research indicates that detectors of machine-generated text are more effective with longer content (Yang et al., 2023b). To leverage this, we aimed to use human samples of maximum possible length. However, the minimum length needed to obtain sufficient samples varied by task. We randomly selected 500 samples of human text from filtered subsets with the following token lengths using Llama2-13B tokenizer (Touvron et al., 2023): 500 tokens for question answering, 400 tokens for code [^4] Length limited to 2500 tokens., 150 tokens for summarization, 275 tokens for dialogue, 500 tokens for reviews, 500 tokens for abstracts, and 500 tokens for translation (Table 3). These 500 samples served as human examples. From them, prompts from the first 100 samples were chosen for use in the generator model, using the input given to the human author as the model prompt. This resulted in a dataset of 500 human examples and 100 machine-generated examples per model for a total of 400 machine-generated examples for each task. This slight data imbalance is intentional to ensure a more accurate TPR@FPR metric because there would likely be more human examples than machine generated examples in practice.
Detection methods show improved performance with longer text sequences (Wu et al., 2023) so we show the statistics of the text in Table 3. Our primary focus was on detectors’ ability to identify AI-generated text while maintaining a low FPR. The longer length of human-generated text is likely to enhance the TPR@FPR by making it easier to detect as human. We considered the AI-generated text sufficiently long for two reasons. First, Li et al. (2024) reports an average AI generation length of 279.99, which is much lower than our average token lengths. Their extensive training and evaluation data support the adequacy of this length for AI content. Second, our models, with a maximum generation length of 512 tokens  [^5] The averages can exceed this number due to different tokenizers and additional tokens to keep text coherent, produced responses indicative of real-world lengths.
Task AI Human
Avg Min Avg Min
Code 486.58 15 4496.88 605
QA 508.01 24 1052.37 501
Summ 410.03 18 191.00 151
Dialogue 380.92 15 402.13 276
Reviews 551.28 24 796.06 501
Abstract 427.92 30 2081.88 501
Translation 525.32 256 772.75 501
Table 3: Average and minimum token counts of machine-generated and human-generated text for each task, tokenized using the Llama2 tokenizer (Touvron et al., 2023). Minimum token counts for human-generated text are omitted as they were previously described.
4.2 Text Generation and Detection Process
Once the prompt samples were selected, we needed to generate positive examples. The process for this can be seen in Figure 1. We employ three different strategies for prompting the models, one is a plain prompt and the other two are adversarial prompts. The first strategy involves using a basic prompt for each domain that explains the goal of the model and the desired output format. The second strategy consists of requesting that the model be as human as possible. The third strategy requests that the model rewrite and improve upon the human written response  [^6] Prompts and templates can be found in the appendix.. The first strategy aims to simulate a basic system prompt that would generally be in place on a model someone is using to generate content. The second strategy simulates the case where a user might try to get the model to generate content that closely resembles human-generated content. The third strategy simulates a scenario where the user writes their own response and simply wants the model to clean it up or make it easier to understand. The outputs of the models were taken as is with no editing. After generating the positive examples, we passed all of the machine-generated and human-generated examples through the detectors. RADAR, Wild, and T5Sentinel all return a percentage probability for each class, and GPTID, Fastdetectgpt, Binoculars, and LogRank return a value representing their score. We do not use any thresholds and take the scores as is for AUROC and TPR@FPR metrics.
Figure 1: Pipeline for prompting and evaluation. Adversarial prompting and rewriting are applied to the LLMs. After collecting machine-generated text, AUROC and TPR@FPR are measured for each detector.
Detector TPR@0.01 TPR@0.05 TPR@0.1 AUROC
Radar 0.05 0.15 0.27 0.6009
Fast-DetectGPT 0.49 0.61 0.68 0.8405
Wild 0.11 0.19 0.29 0.6841
PHD 0.08 0.23 0.37 0.6790
LogRank 0.09 0.40 0.50 0.7763
T5Sentinel 0.03 0.09 0.14 0.5179
Binoculars 0.58 0.67 0.72 0.8485
Table 4: Performance of different detectors across the entire dataset.
5 Results and Analysis
(a) Average TPR@0.01 results for multilingual tasks with normal prompting across all detectors.
(b) Average AUROC results for multilingual tasks with normal prompting across all detectors.
(c) Average TPR@0.01 results for multilingual tasks with template prompting across all detectors.
(d) Average TPR@0.01 results for multilingual tasks with rewrite prompting across all detectors.
Figure 2: Comparison of average AUROC results for multilingual tasks across all detectors using different normal prompting and average TPR@0.01 across all detectors using normal, template, and rewrite prompting. Error bars show maximum and minimum performance across detectors.
(a) Average TPR@0.01 results for English tasks with normal prompting across all detectors.
(b) Average AUROC results for English tasks with normal prompting across all detectors.
(c) Average TPR@0.01 results for English tasks with template prompting across all detectors.
(d) Average TPR@0.01 results for English tasks with rewrite prompting across all detectors.
Figure 3: Comparison of average AUROC results for English tasks across all detectors using different normal prompting and average TPR@0.01 across all detectors using normal, template, and rewrite prompting. Error bars show maximum and minimum performance across detectors.
Table 4 shows the overall performance of each detector across the entire dataset. In this section, we break down the performance of each detector across tasks, languages, and prompt techniques.
5.1 Plain Prompting
We evaluate the AUROC and TPR at 0.01 FPR for machine-generated texts from direct prompting using identical prompts as human written texts. A simple prompt was employed to ensure the generated text was in the correct format and language for the multilingual tasks.
Figures 2(a) and 2(b) show the results for the multilingual tasks and 3(a) and 3(b) show the results for the only English tasks. The results broken down by detector are shown in Appendix A.3. A significant difference is observed in detector performance across languages and tasks, particularly in the multilingual setting as well as across detectors. In the TPR@.01 setting, the difference between the best detector and worst detector is greater than $0.950.95$. Across all detectors we generally see strong results in the English tasks, while the performance drops off in the non-English tasks. In most detectors, in all tasks, they struggle to maintain a strong TPR rate at an FPR rate of 0.01.
For the English-only tasks, most detectors show improved performance in the AUROC, while the TPR@0.01 stays quite low. Despite expectations that the translation domain would be the most challenging due to lower entropy in translated texts, detectors performed reasonably well from the AUROC perspective. The TPR@0.01 graph highlights ongoing challenges in maintaining low false positive rates.
5.2 Adversarial Prompting
Figure 2(c) shows the results on the multilingual tasks where the model was instructed to be "as human as possible." Interestingly, this request had little effect on performance. In the few instances where changes occurred, scores generally increased, suggesting that asking the model to "sound human" may have made its output easier to detect. This aligns with expectations, as large language models are already trained on predominantly human-written texts, and generating more conversational output can make detection more straightforward, as evidenced in dialogue generation tasks.
On the English tasks, as shown in figure, 3(c), the results were similarly unaffected by the human-like request, with some slight score increases where changes were observed. This is especially expected in domains such as reviews, code, and abstracts, which follow specific writing conventions, while tasks like question answering and dialogue generation exhibit more variability and creativity.
Model Code Reviews Abstract QA Summ Dialogue Trans. Arena Score
TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC
GPT-4o 0.02 0.55 0.28 0.63 0.04 0.53 0.05 0.54 0.05 0.50 0.03 0.58 0.03 0.52 1339
Llama-3 0.06 0.56 0.28 0.67 0.21 0.64 0.12 0.60 0.07 0.57 0.08 0.60 0.09 0.56 1152
Mistral 0.02 0.54 0.28 0.65 0.04 0.54 0.10 0.58 0.04 0.51 0.06 0.59 0.05 0.54 1072
Phi-3 0.04 0.57 0.24 0.62 0.13 0.58 0.08 0.58 0.12 0.58 0.13 0.63 0.08 0.56 1066
Table 5: Model performance (AUROC and TPR@0.01) across tasks compared with model generation quality. The Chatbot Arena score is utilized to measure the quality of a model. The higher scores do not correlate with lower detectability of generated content.
5.3 Rewriting
Finally, we show the results for the rewriting prompt for the multilingual tasks in figure 2(d) and for the English tasks in figure 3(d). We observe a notable decrease in TPR@0.01 performance for detectors that previously performed well leading to a drop in the average performance in most tasks. Some of the lower performing did see an increase in performance which is why the average performance in the Code and French Translation tasks are slightly higher. Despite these shifts, the relative performance across tasks remains consistent, indicating an inherent variability in detectability based on the type of task and language.
Figure 4: Correlations between the TPR at various FPR rates and the overall AUROC score. AUROC score is more representative of the middle FPR rates, while this detection task is more concerned with the lower end of FPR.
5.4 TPR@FPR vs AUROC
In this paper, we utilize both the AUROC and TPR@FPR metrics. However, we also argue that TPR at a low FPR is a much more important metric for this detection task. Figure 4 shows the correlation between TPR scores at various FPR rates and the AUROC score for all tasks, detectors, and models used in this research. The AUROC correlates much higher with FPR rates in the 0.4 to 0.6 range and much lower with FPR rates at the edges, less than 0.2 and greater than 0.8. While the 0.75 is still a reasonable correlation value, the AUROC is still much more representative of the middle FPR’s while we are really concerned with the lower FPR’s for this task. This is why we report the TPR@0.01, which is much more representative of the applicability of a detector than the AUROC.
5.5 Output Quality and Detection
Measuring the quality of LLM outputs, especially in creative tasks, remains challenging, making it difficult to determine if higher-quality outputs are harder to detect. Table 5 compares various models’ performance scores and rankings from Chatbot Arena (Chiang et al., 2024), allowing us to explore if output quality affects detectability. The data shows little difference in detectability across models of varying quality, with AUROC and TPR@0.01 scores remaining consistent. This suggests that output quality does not significantly impact the difficulty of detection, though further research is needed for a fuller understanding.
Figure 5: AUROC and TPR@0.01 for each zero-shot method using various underlying models. Only Fast-DetectGPT and Binoculars show a significant change in performance with GPT2-Medium.
5.6 Impact of Model on Zero-shot Methods
Each zero-shot method used in this paper has an underlying model that assists in the detection process. In this paper we consider the model chosen by the respective authors of each detector to be a part of the detector itself. However, we also swapped out each model to directly compare the statistical methods themselves, removing any impact from a specific model on a detector.
Figure 5 shows the results of running each detector across the entire dataset with three different models. There is generally not much of a difference in the ability of a detection method when changing the underlying model. Fast-DetectGPT and Binoculars show a small change in AUROC and a larger change in TPR@0.01 when using the gpt2-medium model (Radford et al., 2019). Gpt2-medium is the oldest model of the three, which likely results in its output logits being different than the generation models more often. This provides some evidence that these zero-shot methods will require updated underlying models to remain successful on more advanced generation models, but more research would need to be conducted.
6 Conclusion
This study evaluates seven advanced detectors across seven tasks and four languages, revealing notable inconsistencies in their detection capabilities. We also examined three different prompting strategies and their impact on detectability, finding that requests for more "human-like" output do not make the text harder to detect, while rewritten human content proves more difficult to identify.
The detection results for both the Translation task and Rewrite prompt are generally lower than the average detectability for other machine generated text. This encourages a discussion about whether this type of text should be detected as machine generated or not. The text may have been machine generated, but it is heavily influenced by human generated text. Specifically in the translation case, the text should match the source text in another language. In the Rewrite case, the model is not generating any new ideas, just improving the readability of the text. It is clear from the results that these cases are harder to detect than when a model has a more open-ended generation. It is likely worth differentiating between these two cases which we leave to future work.
Additionally, this research highlights the limitations of relying on the AUROC metric for assessing machine-generated content detectors. Our findings emphasize the need for robust evaluation methods to develop more reliable detection techniques. The study underscores the challenges in detecting machine-generated text, particularly when human written text was only modified by a language model, and advocates for TPR@FPR as the preferred evaluation metric to better capture detector performance.
7 Limitations
A limitation of this method is the settings in which the human data was collected may vary from the settings in which these detectors will be used. Additionally, some of the datasets we used had collected their data from the internet which raises a concern that some of that data is not completely human generated. This is a challenge that all future detectors will also struggle with when training and evaluating. These results pose the risk of emboldening users to use AI generated content when they otherwise should not because they know detectors cannot be confidently trusted. However, acknowledging this is important to encouraging research into new detection methods and improving current methods.
Acknowledgements
L.L. is partly supported by a gift from Hydrox AI and a CMU CyLab Seed Fund. The views expressed are those of the author and do not reflect the official policy or position of the funding agencies.
Appendix A More Results
This section contains results for detections by models and tasks, the prompts used for plain prompting, and the results by detector.
A.1 Results by Model
The following tables show the results for each detector by generation model and task. As discussed in the paper, there is not a significant difference in detectability of a text by the model that generated that text. More specifically, a higher quality model like GPT-4o is not noticeably harder to detect than a lower quality model like Phi-3. The differences in detectability are more obvious across tasks than generation models.
Model Detector TPR@.01 AUROC
GPT-4o Binoculars 0.14 0.8739
Fast-DetectGPT 0.07 0.8223
LogRank 0.00 0.5722
PHD 0.00 0.3976
Radar 0.00 0.6594
T5Sentinel 0.01 0.4604
Wild 0.00 0.4724
Llama-3 Binoculars 0.45 0.9474
Fast-DetectGPT 0.16 0.8211
LogRank 0.00 0.6060
PHD 0.04 0.4940
Radar 0.05 0.8088
T5Sentinel 0.09 0.5358
Wild 0.06 0.6231
Mistral Binoculars 0.39 0.9681
Fast-DetectGPT 0.16 0.9029
LogRank 0.06 0.4552
PHD 0.06 0.3104
Radar 0.00 0.5787
T5Sentinel 0.02 0.3315
Wild 0.07 0.5412
Phi-3 Binoculars 0.35 0.7825
Fast-DetectGPT 0.21 0.8076
LogRank 0.08 0.5470
PHD 0.14 0.4657
Radar 0.06 0.7700
T5Sentinel 0.03 0.5209
Wild 0.12 0.5779
Table 6: Code
Model Detector TPR@.01 AUROC
GPT-4o Binoculars 0.47 0.8676
Fast-DetectGPT 0.35 0.8918
LogRank 0.01 0.5972
PHD 0.01 0.4758
Radar 0.02 0.3474
T5Sentinel 0.01 0.4909
Wild 0.02 0.5088
Llama-3 Binoculars 0.66 0.9501
Fast-DetectGPT 0.68 0.9465
LogRank 0.03 0.7816
PHD 0.02 0.6637
Radar 0.13 0.6581
T5Sentinel 0.03 0.4995
Wild 0.06 0.6024
Mistral Binoculars 0.56 0.8954
Fast-DetectGPT 0.57 0.8852
LogRank 0.04 0.7052
PHD 0.04 0.6261
Radar 0.04 0.5776
T5Sentinel 0.02 0.5078
Wild 0.04 0.5664
Phi-3 Binoculars 0.48 0.8188
Fast-DetectGPT 0.41 0.8424
LogRank 0.13 0.7448
PHD 0.13 0.6553
Radar 0.07 0.5966
T5Sentinel 0.02 0.5787
Wild 0.07 0.5923
Table 7: Question Answering
Model Detector TPR@.01 AUROC
GPT-4o Binoculars 0.05 0.6533
Fast-DetectGPT 0.11 0.6981
LogRank 0.23 0.7070
PHD 0.00 0.4922
Radar 0.00 0.2085
T5Sentinel 0.01 0.4339
Wild 0.11 0.4753
Llama-3 Binoculars 0.33 0.8134
Fast-DetectGPT 0.19 0.7540
LogRank 0.61 0.8941
PHD 0.30 0.6642
Radar 0.08 0.6605
T5Sentinel 0.05 0.4720
Wild 0.15 0.6570
Mistral Binoculars 0.06 0.6011
Fast-DetectGPT 0.08 0.5862
LogRank 0.30 0.7635
PHD 0.01 0.5216
Radar 0.00 0.3428
T5Sentinel 0.02 0.4256
Wild 0.13 0.5931
Phi-3 Binoculars 0.48 0.7501
Fast-DetectGPT 0.30 0.7135
LogRank 0.60 0.9049
PHD 0.33 0.7668
Radar 0.09 0.7492
T5Sentinel 0.01 0.3561
Wild 0.38 0.8890
Table 8: Summarization
Model Detector TPR@.01 AUROC
GPT-4o Binoculars 0.68 0.9362
Fast-DetectGPT 0.52 0.9277
LogRank 0.37 0.7857
PHD 0.07 0.6279
Radar 0.03 0.6143
T5Sentinel 0.06 0.5045
Wild 0.01 0.5633
Llama-3 Binoculars 0.80 0.9695
Fast-DetectGPT 0.67 0.9487
LogRank 0.42 0.8235
PHD 0.24 0.6772
Radar 0.14 0.6907
T5Sentinel 0.04 0.5265
Wild 0.03 0.6378
Mistral Binoculars 0.64 0.9150
Fast-DetectGPT 0.56 0.9246
LogRank 0.33 0.7827
PHD 0.17 0.6514
Radar 0.19 0.6989
T5Sentinel 0.02 0.4949
Wild 0.02 0.6044
Phi-3 Binoculars 0.69 0.9243
Fast-DetectGPT 0.59 0.8506
LogRank 0.60 0.9206
PHD 0.37 0.7745
Radar 0.04 0.7626
T5Sentinel 0.11 0.6134
Wild 0.10 0.6505
Table 9: Dialogue
Model Detector TPR@.01 AUROC
GPT-4o Binoculars 0.39 0.8584
Fast-DetectGPT 0.33 0.8895
LogRank 0.00 0.6474
PHD 0.00 0.2399
Radar 0.01 0.1890
T5Sentinel 0.00 0.3086
Wild 0.06 0.6178
Llama-3 Binoculars 0.71 0.9248
Fast-DetectGPT 0.70 0.9352
LogRank 0.12 0.8505
PHD 0.09 0.5909
Radar 0.29 0.6430
T5Sentinel 0.16 0.6843
Wild 0.42 0.8185
Mistral Binoculars 0.44 0.8672
Fast-DetectGPT 0.38 0.8937
LogRank 0.01 0.6964
PHD 0.01 0.3262
Radar 0.03 0.1770
T5Sentinel 0.05 0.4642
Wild 0.04 0.5601
Phi-3 Binoculars 0.38 0.5372
Fast-DetectGPT 0.43 0.7147
LogRank 0.30 0.8252
PHD 0.48 0.7344
Radar 0.42 0.8498
T5Sentinel 0.01 0.3748
Wild 0.49 0.9189
Table 10: Abstract
Model Detector TPR@.01 AUROC
GPT-4o Binoculars 0.69 0.9247
Fast-DetectGPT 0.65 0.8847
LogRank 0.31 0.8778
PHD 0.00 0.7663
Radar 0.12 0.8791
T5Sentinel 0.02 0.5718
Wild 0.00 0.9249
Llama-3 Binoculars 0.80 0.9662
Fast-DetectGPT 0.72 0.9305
LogRank 0.65 0.9240
PHD 0.47 0.8589
Radar 0.39 0.8883
T5Sentinel 0.11 0.5814
Wild 0.30 0.9541
Mistral Binoculars 0.79 0.9659
Fast-DetectGPT 0.74 0.9358
LogRank 0.60 0.9249
PHD 0.06 0.8347
Radar 0.37 0.9015
T5Sentinel 0.07 0.6376
Wild 0.42 0.9574
Phi-3 Binoculars 0.70 0.8889
Fast-DetectGPT 0.49 0.8178
LogRank 0.40 0.8272
PHD 0.17 0.8137
Radar 0.47 0.9024
T5Sentinel 0.03 0.4166
Wild 0.25 0.9410
Table 11: Reviews
Model Detector TPR@.01 AUROC
GPT-4o Binoculars 0.12 0.7020
Fast-DetectGPT 0.05 0.6539
LogRank 0.02 0.6059
PHD 0.02 0.5611
Radar 0.01 0.6624
T5Sentinel 0.01 0.3732
Wild 0.13 0.5788
Llama-3 Binoculars 0.54 0.8739
Fast-DetectGPT 0.40 0.8146
LogRank 0.23 0.7281
PHD 0.08 0.6874
Radar 0.18 0.8910
T5Sentinel 0.06 0.4902
Wild 0.37 0.7462
Mistral Binoculars 0.31 0.7781
Fast-DetectGPT 0.14 0.7139
LogRank 0.05 0.6167
PHD 0.03 0.6115
Radar 0.08 0.8602
T5Sentinel 0.02 0.3869
Wild 0.19 0.6562
Phi-3 Binoculars 0.36 0.7671
Fast-DetectGPT 0.20 0.6446
LogRank 0.09 0.6377
PHD 0.30 0.7130
Radar 0.36 0.9726
T5Sentinel 0.01 0.4116
Wild 0.45 0.8199
Table 12: Translation
A.2 Plain Prompts
Table 13 shows the prompts used for each task in the plain prompting. The prompts we used were intentionally very simple and not overly instructive. This is because we wanted to replicate a realistic scenario of an average person prompting a language model. We performed small ablations on these prompts and found no difference in detectability.
Task Prompt
Code You are a helpful code assistant that can teach a junior developer how to code. Your language of choice is Python. Don’t explain the code, just generate the code block itself.
Question Answering You are a helpful question answering assistant who will answer a single question as completely as possible given the information in the question. Do NOT use any markdown, bullet, or numbered list formatting. The assistant will use ONLY paragraph formatting. **Respond only in {language}**
Summarization You are a helpful summarization assistant that will summarize a given article. Provide only the summarization in paragraph formatting. Do not introduce the summary. **Respond in {language}**
Dialogue You are a helpful dialogue generation assistant that will generate a dialogue between people given a short paragraph describing the people involved. Provide only the dialogue. Do not introduce the dialogue. **Respond in {language}**
Abstract Writing You are a helpful abstract writing assistant. You will write an abstract given the content of a paper. Do not provide any other text. You will only provide an abstract.
Review Writing You are a helpful conference paper review assistant. Please provide a detailed review of the following paper, including its strengths, weaknesses, and suggestions for improvement.
Translation You are a helpful translation assistant that will translate a given text into English. Provide only the translation and nothing else.
Rewriting You are a helpful writing assistant. Rewrite the following text to improve clarity and professionalism. Do not provide any other text. Only provide the rewritten text.
Table 13: The table shows the prompts used in the plain prompting. For GPT, these were used as system prompts, and for huggingface models they were prepended to the questions.
A.3 Results by Detector
This section shows the numerical value of each detector on each task. In the paper we display graphs representing most of these values but show all of the numbers here for reference. The TPR@.01 and AUROC change significantly across tasks for every detector signifying that these detectors are not equally capable of detecting all types of machine generated text.
Code Reviews Abstract Translation ES Translation FR Translation ZH
TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC TPR AUC
Radar 0.0258 0.7042 0.3358 0.8928 0.1858 0.4647 0.1806 0.8528 0.1475 0.8565 0.3625 0.8307
Fast-DetectGPT 0.1508 0.8385 0.6475 0.8922 0.4608 0.8583 0.1782 0.6337 0.0817 0.6078 0.3333 0.7953
Wild 0.0608 0.5537 0.2425 0.9443 0.2517 0.7288 0.1759 0.7433 0.0875 0.5455 0.5358 0.9049
PHD 0.0625 0.4169 0.1758 0.8184 0.1467 0.4729 0.0556 0.5264 0.0075 0.6328 0.2283 0.7915
LogRank 0.0367 0.5451 0.4883 0.8884 0.1075 0.7549 0.0509 0.4446 0.0600 0.5900 0.1792 0.7820
T5Sentinel 0.0400 0.4621 0.0575 0.5519 0.0575 0.4580 0.0231 0.2582 0.0025 0.3473 0.0767 0.5286
Binoculars 0.3317 0.8930 0.7450 0.9364 0.4783 0.7969 0.1944 0.6989 0.1492 0.6754 0.5233 0.8807
Table 14: Detector performance (AUROC and TPR@0.01) across tasks.
QA EN QA ES QA FR QA ZH
TPR AUC TPR AUC TPR AUC TPR AUC
Radar 0.1225 0.7542 0.0100 0.4832 0.0525 0.5008 0.0400 0.3730
Fast-DetectGPT 0.5450 0.9063 0.4233 0.8522 0.6483 0.9390 0.5808 0.8864
Wild 0.1375 0.7437 0.0308 0.5019 0.0183 0.4660 0.0275 0.6180
PHD 0.0600 0.4694 0.0792 0.6228 0.0158 0.6371 0.0733 0.7138
LogRank 0.0725 0.7517 0.0725 0.6427 0.0200 0.7741 0.1308 0.8323
T5Sentinel 0.0558 0.6007 0.0075 0.4353 0.0025 0.4131 0.0175 0.6688
Binoculars 0.6950 0.9271 0.5292 0.8295 0.7250 0.9326 0.4908 0.8785
Table 15: Detector performance (AUROC and TPR@0.01) across multilingual QA tasks.
Summ EN Summ ES Summ FR Summ ZH
TPR AUC TPR AUC TPR AUC TPR AUC
Radar 0.0825 0.5721 0.0125 0.3839 0.0208 0.3966 0.2433 0.6940
Fast-DetectGPT 0.2175 0.8185 0.1408 0.6446 0.1058 0.6169 0.1950 0.6729
Wild 0.1250 0.7436 0.2683 0.6027 0.2350 0.5578 0.4392 0.8499
PHD 0.1292 0.5907 0.1417 0.6107 0.1617 0.5875 0.1933 0.6631
LogRank 0.7517 0.9705 0.4425 0.8902 0.4217 0.8754 0.4533 0.8317
T5Sentinel 0.1333 0.6275 0.0042 0.2728 0.0208 0.3219 0.0183 0.3828
Binoculars 0.3792 0.7916 0.2225 0.6866 0.1942 0.6482 0.1333 0.6935
Table 16: Detector performance (AUROC and TPR@0.01) across multilingual summarization tasks.
Dialogue EN Dialogue ES Dialogue FR Dialogue ZH
TPR AUC TPR AUC TPR AUC TPR AUC
Radar 0.7175 0.9344 0.0583 0.6827 0.0417 0.5549 0.1317 0.6120
Fast-DetectGPT 0.7475 0.9508 0.4300 0.8847 0.6158 0.9098 0.4217 0.9011
Wild 0.0292 0.8720 0.0775 0.6266 0.1000 0.6035 0.0617 0.3290
PHD 0.0425 0.6224 0.1892 0.7168 0.1842 0.7401 0.3835 0.7962
LogRank 0.8583 0.9889 0.4250 0.8824 0.4283 0.8987 0.5208 0.8870
T5Sentinel 0.1708 0.6628 0.0467 0.6010 0.0508 0.4639 0.0217 0.3446
Binoculars 0.7767 0.9489 0.7700 0.9407 0.7125 0.9393 0.6467 0.9475
Table 17: Detector performance (AUROC and TPR@0.01) across multilingual dialogue tasks.